2w(5w+3)=0 solve using zero-factor property
Answer by stanbon(75887)

2w(5w+3)=0 solve using zero-factor property
----
2w = 0 or 5w+3 = 0
---
w = 0 or w = -3/5
======================
Cheers,
Stan H.
======================